Problem 13-2 EBIT, Taxes, and Leverage [LO 2] Kaelea, Inc., has no debt outstanding and a total market value of $100,000. Earnings before interest and taxes, EBIT, are projected to be $8,400 if economic conditions are normal. If there is strong expansion in the economy, then EBIT will be 24 percent higher. If there is a recession, then EBIT will be 31 percent lower. The company is considering a $35,000 debt issue with an interest rate of 6 percent. The proceeds will be used to repurchase shares of stock. There are currently 4,000 shares outstanding. Assume the company has a tax rate of 35 percent.

a. Calculate earnings per share, EPS, under each of the three economic scenarios before any debt is issued. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) EPS Recession $ Normal $ Expansion $

b. Calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to the nearest whole number, e.g., 32.) %ΔEPS Recession % Expansion % Assume the company goes through with recapitalization.

c. Calculate earnings per share, EPS, under each of the three economic scenarios after the recapitalization. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) EPS Recession $ Normal $ Expansion $

d. Calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) %ΔEPS Recession % Expansion %
Answer #1
a
EPS = EBIT*(1-tax rate)/shares outstanding
Recession
EPS = EBIT*(1-recession impact%)*(1-tax rate)/shares outstanding
EPS=8400*(1-0.31)*(1-0.35)/4000
EPS=0.94
Normal
EPS = EBIT*(1-tax rate)/shares outstanding
EPS=8400*(1-0.35)/4000
EPS=1.37
Expansion
EPS = EBIT*(1+Growth impact%)*(1-tax rate)/shares outstanding
EPS=8400*(1+0.24)*(1-0.35)/4000
EPS=1.69
b
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(0.9419/1.365-1)*100
=-31%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(1.6926/1.365-1)*100
=24%
c
New no. of shares = old shares-debt/(Market value/old shares)
=4000-35000/(100000/4000)
=2600
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
Recession
EPS = (EBIT*(1-recession impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(8400*(1-0.31)-35000*0.06)*(1-0.35)/2600
EPS=0.92
Normal
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
EPS=(8400-35000*0.06)*(1-0.35)/2600
EPS=1.58
Expansion
EPS = (EBIT*(1+growth impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(8400*(1+0.24)-35000*0.06)*(1-0.35)/2600
EPS=2.08
d
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(0.924/1.575-1)*100
=-41%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(2.079/1.575-1)*100
=32%